OpenStudy (anonymous):

help me write the decimal as a fraction or a mixed number in simplest form. 3.4 with a line over the 4.

OpenStudy (anonymous):

.75

OpenStudy (anonymous):

that doesn't even make sense, it says turn the decimal into a fraction or mixed number.

OpenStudy (anonymous):

1/9=0.11111111111111111111 forever 2/9=0.222222222222222222222 forever 4/9=0.44444444444444444 forever